# Computing the Intrinsic Delaunay Triangulation of a Closed Polyhedral Surface

### v1

@@ 4.2 Deleting vertices
Appendix B details how to modify the portalgon $T$ to perform DeleteVertices. Afterward the polygons of $T$ are usually not triangles anymore, but this will be solved by applying InsertEdges after each application of DeleteVertices. Observe that in DeleteVertices we delete only vertices of degree smaller than or equal to six. Informally, the reason is that deleting a weak vertex of degree $d\geq 3$ creates a face of degree $d$ around it. We then insert d3d-3 edges in this face when applying InsertEdges. The problem is that only a constant number of edges can be inserted in each face without risking to destroy our improvements on the geometry of the tessellation. This is why we make sure that $d=O(1)$ beforehand. The exact bound on $d$ is not really important (although changing it would change some constants of the algorithm), but it must be at least six so that we can still remove a fraction of the excess vertices this way, at least when most of them are strong. Similar ideas can be found in the literature, see for example Kirkpatrick [kirkpatrick1983optimal, Lemma 3.2].
@@ 4.4 Data structure for marking bifaces as inactive
Figure 3: Data structure for Algorithm: a portalgon whose polygons are partitioned, here by color, inducing sub-portalgons called regions, and a region singularized as active, here in red.
See Figure 3. The data structure maintains a portalgon $R$ together with a partition of the polygons of $R$. Each set $X$ of polygons in the partition defines a sub-portalgon of $R$ which we call region. One region is singularized as the active region $R_{A}$. The other regions are inactive. Note that the surface of the active region may be disconnected, and that the surfaces of distinct inactive regions may be adjacent.
@@ 4.5 Algorithm
Algorithm. Given a triangular portalgon $T$ whose surface $\mathcal{S}(T)$ is flat, and $N\geq 1$, do the following. Initialize the data structure by letting $R$ be the input portalgon $T$, and the active region $R_{A}$ be $R$ itself, without inactive region. Repeat $N$ times the following:
@@
      Apply Gardening. Then apply SimplifyTubes to $R_{A}$ but in step 2 whenever the good biface $B$ is thin mark $B$ as inactive. Apply Gardening again.
@@ 5.1 Combinatorial analysis
The iteration starts with InsertVertices. This is the only moment where $m_{A}$ may increase, and we prove that $m_{A}$ is multiplied by at most a constant factor. Then the iteration applies the interior loop, and we claim that, as long as $m_{A}$ exceeds $n$ by a constant factor, $m_{A}$ is divided by at least a constant factor by each iteration of the interior loop. We show that this claim implies the lemma as the interior loop is applied sufficiently many times to counteract the initial increase of $m_{A}$. To prove the claim, we show that for DeleteVertices to remove a fraction of the vertices of $R_{A}^{1}$, it suffices that $m_{A}$ vastly exceeds the genus and the number of boundary components of $\mathcal{S}(R_{A})$, and that almost all of the vertices of $R_{A}^{1}$ are weak. We show that this is ensured by first applying Gardening and SimplifyTubes.
@@ 5.2 Enclosure
To analyze Algorithm from a geometric point of view, we introduce, on the segments of a flat surface $S$, a parameter that we call enclosure. So consider a segment $e$ of $S$. See Figure 4.
@@
Figure 4: The red loop encloses the blue segment in the surface.
@@ 5.3.1 The separating loops are not very enclosed
Proof 5.6 (Proof of Lemma 5.4).
Only step 2 of SimplifyTubes may create a separating loop, by marking a thin biface $B$ as inactive. Then $B$ is is never touched again by the algorithm. So the algorithm maintains the invariant that every separating loop $e$ is adjacent to the surface of at least one inactive region that is a thin biface. So $c_{S}(e)\leq 2$ by Proposition 5.5.
@@ 5.3.2 The very enclosed edges shorten exponentially fast
Proof 5.10 (Proof of Lemma 5.8).
First observe that $e^{\prime}$ is not included in the boundary of $\mathcal{S}(R_{A}^{\prime})$ because $e^{\prime}$ is enclosed and thus not included in the boundary of $S$, and because $e^{\prime}$ is not a separating loop by Lemma 5.4. So there is an edge $e$ of $R_{A}^{1}$ such that $e^{\prime}$ is one of the two half-segments obtained after the insertion of the middle point of $e$ as a vertex. Then $\ell(e)=2\ell(e^{\prime})$. And $c_{S}(e)\geq c_{S}(e^{\prime})$ by Lemma 5.9.
@@
Proof 5.13 (Proof of Lemma 5.11).
Here we crucially use the fact that every polygon of $R_{A}$ has degree at most six. so that at most three edges are inserted within the polygon. Indeed either $e^{\prime}$ was already an edge of $R_{A}^{1}$ and there is nothing to do, or $e^{\prime}$ has been inserted in some face $F$ of $R_{A}^{1}$. At most three edges were inserted in $F$, and Proposition 5.12 applied at most three times gives a boundary edge $e$ of $F$ such that $c_{S}(e)\geq c_{S}(e^{\prime})-12$ and $\ell(e)\geq(1-12/c_{S}(e^{\prime}))\ell(e^{\prime})$.
@@
Proof 5.15 (Proof of Proposition 5.7).
Consider the active regions $R_{A}$ and $R^{\prime}_{A}$ respectively at the beginning of the algorithm, and after $i$ iterations of the main loop. Assume that there is an edge $e^{\prime}$ in $R_{A}^{\prime 1}$ such that $c_{S}(e^{\prime})>22000\cdot i$. During those $i$ iterations there has been $i$ applications of InsertVertices, $351i$ applications of InsertEdges, and $350i$ applications of SimplifyTubes. Also $12\cdot 351i+5\cdot 350i<11000i$. So Lemma 5.8, Lemma 5.11, and Lemma 5.14 imply that there is an edge $e$ in $R_{A}^{1}$ such that $\ell(e)\geq 2^{i}(1-11000i/c_{S}(e^{\prime}))\ell(e^{\prime})>2^{i-1}\ell(e^{\prime})$. And $\ell(e)\leq L$ because $e$ belongs to the input triangulation $T^{1}$.
@@
Our second claim is that in the end every edge $e$ of $R^{1}$ satisfies $h_{S}(e)=O(\log(n)\cdot\log^{2}(2+L/s))$. Let us prove this second claim. First observe that if $e$ is in $R_{A}^{1}$ then $c_{S}(e)<22000\log(2+L/s)$, for otherwise Proposition 5.7 would imply $\ell(e)<2s$, implying that no loop encloses $e$ in $S$, a contradiction. In this case $h_{S}(e)=O(\log(2+L/s)\cdot(\log(n)+\log(2+L/s)))$ by Proposition 5.3, and we are done. Every other edge of $R^{1}$ belongs to the 1-skeleton of an inactive good biface $B$. Every boundary edge $e$ of $B^{1}$ is either a boundary component of $S$ or a separating loop, so $c_{S}(e)\leq 2$ by Lemma 5.4, and so $h_{S}(e)=O(\log(n)+\log(2+L/s))$ by Proposition 5.3. Every interior edge $f$ of $B^{1}$ then satisfies $h_{S}(f)=O(\log(n)+\log(2+L/s))$ by Lemma 3.1. This proves the second claim, and the proposition.
Acknowledgments.
The authors thanks Éric Colin de Verdière, Vincent Despré, and anonymous reviewers, whose numerous suggestions improved the presentation of the paper.
@@ Appendix A Appendix of Section 3
Proof A.1 (Proof of Lemma 3.1).
Among the two interior edges of $B^{1}$, let $f$ be a shortest one. Let $g\neq f$ be the other interior edge of $B^{1}$. Let $p$ be a shortest path in $\mathcal{S}(B)$. The relative interior $\mathring{p}$ of $p$ cannot intersect the relative interior of $f$ twice for those intersections would be crossings and $p$ and $f$ are both shortest paths because $B$ is good. So $\mathring{p}$ intersects $f$ less than four times. Then $\mathring{p}$ cannot intersect the relative interior of $g$ five times, for those intersections would be crossings, and $\mathring{p}$ would intersect $f$ in-between any two consecutive crossings with the relative interior of $g$. Altogether $p$ intersects $f$ and $g$ at most six times each.
@@
Proposition 3.2 is similar to but different from a result of Löffler, Ophelders, Silveira, and Staals [portalgons, Theorem 45] (building upon a ray shooting algorithm of Erickson and Nayyeri [erickson2012tracing]), in which the authors provide an algorithm to transform a biface into a portalgon of bounded happiness, and of bounded combinatorial complexity. They extend their result from bifaces to portalgons $X$ such that the dual graph of $X^{1}$ in $\mathcal{S}(X)$ has at most one simple cycle, but unfortunately this does not include tubes. We extend their result to tubes to prove Proposition 3.2, reusing some of ideas developed in the core of the paper.
We need a few lemmas. The following is a corollary of [portalgons, Theorem 45]:
@@
Proof A.3.
By the result of Löffler, Ophelders, Silveira, and Staals [portalgons, Theorem 45] we can compute in $O(1+\log h)$ time a portalgon $T$, whose surface is $\mathcal{S}(B)$, whose happiness is $O(1)$, and whose 1-skeleton $T^{1}$ has $O(1)$ edges. Without loss of generality the two vertices $b_{0}$ and $b_{1}$ of $B^{1}$ are also vertices of $T^{1}$, and we know which vertices of the polygons of $T$ correspond to $b_{0}$ and $b_{1}$.
We now describe how to compute, in constant time, from $T$, a good biface of $\mathcal{S}(T)$. The key thing is that we can exploit the fact that $T$ has $O(1)$ combinatorial complexity and happiness to compute by exhaustive search. First compute, in constant time, by exhaustive search, a shortest path $q$ between $b_{0}$ and $b_{1}$ in $\mathcal{S}(T)$: represent $q$ by its pre-image in the polygons of $T$. Then cut the polygons of $T$ along the pre-image of $q$: every time a polygon is cut in two along a segment $a$, the two edges issued of $a$ are not matched in the resulting portalgon (the goal is to cut the surface of $T$, not just changing $T$). Consider the resulting portalgon $D$. Then $\mathcal{S}(D)$ is homeomorphic to a closed disk. The two endpoints $b_{0}$ and $b_{1}$ of $q$ become a set $V$ of four vertices of $D^{1}$ that lie on the boundary of $\mathcal{S}(D)$. Every singularity of $\mathcal{S}(D)$ lies on the boundary of $\mathcal{S}(D)$ and belongs to $V$. Now replace $D$ by a triangular portalgon $D^{\prime}$, of the same surface, and such that the vertex set of $D^{\prime 1}$ is exactly $V$. This can be done for example by iteratively inserting vertex-to-vertex arcs in the faces of $D^{1}$ to make $D^{1}$ a triangulation, and by deleting a vertex $v$ of $D^{1}$ and its incident edges. When $v$ lies on the boundary of $\mathcal{S}(D)$, only the edges whose relative interior is included in the interior of $\mathcal{S}(D)$ are deleted. In the end, identify back the occurrences of $q$ on the boundary of $\mathcal{S}(D^{\prime})$, by matching the two corresponding sides of polygons in $D^{\prime}$, thereby obtaining a biface $B^{\prime}$ of $\mathcal{S}(B)$ such that $q$ is an interior edge of $B^{\prime}$. Change the other interior edge of $B^{\prime}$ if necessary so that $B^{\prime}$ is good.
@@
Using the claim immediately the intersection of $p$ and $e$ has $O(1)$ connected components, so $p$ writes as a concatenation of $k=O(1)$ paths $p_{1},\dots,p_{k}$ such that for every $1\leq i\leq k$ the path $p_{i}$ is either included in $e$ or its relative interior is disjoint from $e$. Every edge $f\neq e$ of $T^{1}$ intersects $p_{i}$ less than 7 times: if $f$ is included in the boundary of $\mathcal{S}(T)$ then $f$ intersects $p_{i}$ at most once, otherwise Lemma 3.1 applies. So $f$ intersects $p$ less than $O(1)$ times. We proved that the segment-happiness of $T$ is $O(1)$. Then the happiness of $T$ is also $O(1)$ because the polygons of $T$ are all triangles. So we can compute a good biface whose surface is that of $T$ in constant time, exactly as in the proof of Lemma A.2.
@@
Proof A.7.
We may assume without loss of generality that no vertex of $Y$ in the interior of $S$ is incident to a loop edge, by cutting $S$ open at an interior loop edge and recursing on the resulting two triangulations otherwise. Euler’s formula gives $m-m_{1}+m_{2}=0$, where $m$, $m_{1}$, and $m_{2}$ count respectively the vertices, edges, and faces of $Y^{1}$. Double counting gives $3m_{2}=2m_{1}-2$ and $\sum_{v}\deg v=2m_{1}$, where the sum is over the vertices $v$ of $Y$. Then $\sum_{v}(6-\deg v)=4$. The two vertices of $Y$ on the boundary of $S$ have degree greater than or equal to four. So in the interior of $S$ every vertex of degree greater than ten must be compensated by a vertex of degree smaller than or equal to ten.
@@
Let us now prove the claim. Consider the number $m^{\prime}$ of vertices of $X^{1}$ not incident to any loop edge that lie in the interior of $\mathcal{S}(X)$. By Lemma A.6, if $m^{\prime}>0$ before an iteration of the algorithm, then at least $m^{\prime}/2$ such vertices have degree smaller than or equal to ten. So $V$ contains at least $m^{\prime}/14$ vertices, which are deleted. Every non-deleted vertex that was incident to a loop edge before the iteration remains incident to a loop edge after the iteration. We proved that $m^{\prime}$ is divided by at least $14/13$ during the iteration, which proves the claim.
Proof A.10 (Proof of Proposition 3.2).
Apply Lemma A.8, and replace $X$ in $O(n\log n)$ time by a concatenation of less than $3n$ bifaces whose edges are smaller than $(3n)^{c}L$ for some constant $c>0$. Each biface $B$ has segment-happiness $O(1+(3n)^{c}L/s)$; indeed the systole of $\mathcal{S}(B)$ is greater than or equal to the systole of $X$, so every segment $e$ in $\mathcal{S}(B)$ satisfies $h_{\mathcal{S}(B)}(e)=O(1+\ell(e)/s)$. Replace $B$ by a good biface whose surface is that of $B$ in $O(\log(n)+\log(2+L/s))$ time with Lemma A.2. Doing so for all bifaces takes $O(n\cdot(\log(n)+\log(2+L/s)))$ time in total. We crudely bound this running time from above by $O(n\log(n)\cdot\log(2+L/s))$. In the end apply Lemma A.4 repeatedly to merge those $O(n)$ good bifaces into a single good biface, in $O(n)$ total time.

### v2

@@ 4.2 Deleting vertices
Appendix B details how to modify the portalgon $T$ to perform DeleteVertices. Afterward the polygons of $T$ are usually not triangles anymore, but this will be solved by applying InsertEdges after each application of DeleteVertices. Observe that in DeleteVertices we delete only vertices of degree smaller than or equal to six. Informally, the reason is that deleting a weak vertex of degree $d\geq 3$ creates a face of degree $d$ around it. We then insert d3d-3 edges in this face when applying InsertEdges. The problem is that only a constant number of edges can be inserted in each face without risking to destroy our improvements on the geometry of the tessellation. This is why we make sure that $d=O(1)$ beforehand. The exact bound on $d$ is not really important (although changing it would change some constants of the algorithm), but it must be at least six so that we can still remove a fraction of the excess vertices this way, at least when most of them are strong. Similar ideas can be found in the literature, see for example Kirkpatrick [18, Lemma 3.2].
@@ 4.4 Data structure for marking bifaces as inactive
Figure 4.1: Data structure for Algorithm: a portalgon whose polygons are partitioned, here by color, inducing sub-portalgons called regions, and a region singularized as active, here in red.
See Figure 4.1. The data structure maintains a portalgon $R$ together with a partition of the polygons of $R$. Each set $X$ of polygons in the partition defines a sub-portalgon of $R$ which we call region. One region is singularized as the active region $R_{A}$. The other regions are inactive. Note that the surface of the active region may be disconnected, and that the surfaces of distinct inactive regions may be adjacent.
@@ 4.5 Algorithm
Algorithm. Given a triangular portalgon $T$ whose surface $\mathcal{S}(T)$ is flat, and $N\geq 1$, do the following. Initialize the data structure by letting $R$ be the input portalgon $T$, and by letting the active region $R_{A}$ be $R$ itself, without inactive region. Repeat $N$ times the following:
@@
      Apply Gardening. Then apply SimplifyTubes to $R_{A}$ but in step 2 of SimplifyTubes, whenever $B$ is thin, mark $B$ as inactive. Apply Gardening again.
@@ 5.1 Combinatorial analysis
The iteration starts with InsertVertices. This is the only moment where $m_{A}$ may increase, and we prove that $m_{A}$ is multiplied by at most a constant factor. Then the iteration applies the interior loop, and we claim that, as long as $m_{A}$ exceeds $n$ by a constant factor, $m_{A}$ is divided by at least a constant factor by each iteration of the interior loop. We show that this claim implies the lemma as the interior loop is applied sufficiently many times to counteract the initial increase of $m_{A}$. To prove the claim, we show that for DeleteVertices to remove a fraction of the vertices of $R_{A}^{1}$, it suffices that $m_{A}$ vastly exceeds the genus and the number of boundary components of $\mathcal{S}(R_{A})$, and that almost all of the vertices of $R_{A}^{1}$ are weak. We show that this is ensured by first applying Gardening and SimplifyTubes. ∎
@@ 5.2 Enclosure
To analyze Algorithm from a geometric point of view, we introduce, on the segments of a flat surface $S$, a parameter that we call enclosure. So consider a segment $e$ of $S$. See Figure 5.1.
@@
Figure 5.1: The red loop encloses the blue segment in the surface.
@@ 5.3.1 The separating loops are not very enclosed
Proof of Lemma 5.3.
Only step 2 of SimplifyTubes may create a separating loop, by marking a thin biface $B$ as inactive. Then $B$ is is never touched again by the algorithm. So the algorithm maintains the invariant that every separating loop $e$ is adjacent to the surface of at least one inactive region that is a thin biface. So $c_{S}(e)\leq 2$ by Proposition 5.4. ∎
@@ 5.3.2 The very enclosed edges shorten exponentially fast
Proof of Lemma 5.6.
First observe that $e^{\prime}$ is not included in the boundary of $\mathcal{S}(R_{A}^{\prime})$ because $e^{\prime}$ is enclosed and thus not included in the boundary of $S$, and because $e^{\prime}$ is not a separating loop by Lemma 5.3. So there is an edge $e$ of $R_{A}^{1}$ such that $e^{\prime}$ is one of the two half-segments obtained after the insertion of the middle point of $e$ as a vertex. Then $\ell(e)=2\ell(e^{\prime})$. And $c_{S}(e)\geq c_{S}(e^{\prime})$ by Lemma 5.7. ∎
@@
Proof of Lemma 5.8.
Here we crucially use the fact that every polygon of $R_{A}$ has degree at most six. so that at most three edges are inserted within the polygon. Indeed either $e^{\prime}$ was already an edge of $R_{A}^{1}$ and there is nothing to do, or $e^{\prime}$ has been inserted in some face $F$ of $R_{A}^{1}$. At most three edges were inserted in $F$, and Proposition 5.9 applied at most three times gives a boundary edge $e$ of $F$ such that $c_{S}(e)\geq c_{S}(e^{\prime})-12$ and $\ell(e)\geq(1-12/c_{S}(e^{\prime}))\ell(e^{\prime})$. ∎
@@
Proof of Proposition 5.5.
Consider the active regions $R_{A}$ and $R^{\prime}_{A}$ respectively at the beginning of the algorithm, and after $i$ iterations of the main loop. Assume that there is an edge $e^{\prime}$ in $R_{A}^{\prime 1}$ such that $c_{S}(e^{\prime})>22000\cdot i$. During those $i$ iterations there has been $i$ applications of InsertVertices, $351i$ applications of InsertEdges, and $350i$ applications of SimplifyTubes. Also $12\cdot 351i+5\cdot 350i<11000i$. So Lemma 5.6, Lemma 5.8, and Lemma 5.10 imply that there is an edge $e$ in $R_{A}^{1}$ such that $\ell(e)\geq 2^{i}(1-11000i/c_{S}(e^{\prime}))\ell(e^{\prime})>2^{i-1}\ell(e^{\prime})$. And $\ell(e)\leq L$ because $e$ belongs to the input triangulation $T^{1}$. ∎
@@
Our second claim is that in the end every edge $e$ of $R^{1}$ satisfies $h_{S}(e)=O(\log(n)\cdot\log^{2}(2+L/s))$. Let us prove this second claim. First observe that if $e$ is in $R_{A}^{1}$ then $c_{S}(e)<22000\log(2+L/s)$, for otherwise Proposition 5.5 would imply $\ell(e)<2s$, implying that no loop encloses $e$ in $S$, a contradiction. In this case $h_{S}(e)=O(\log(2+L/s)\cdot(\log(n)+\log(2+L/s)))$ by Proposition 5.2, and we are done. Every other edge of $R^{1}$ belongs to the 1-skeleton of an inactive good biface $B$. Every boundary edge $e$ of $B^{1}$ is either a boundary component of $S$ or a separating loop, so $c_{S}(e)\leq 2$ by Lemma 5.3, and so $h_{S}(e)=O(\log(n)+\log(2+L/s))$ by Proposition 5.2. Every interior edge $f$ of $B^{1}$ then satisfies $h_{S}(f)=O(\log(n)+\log(2+L/s))$ by Lemma 3.1. This proves the second claim, and the proposition. ∎
@@ Appendix A Appendix of Section 3
Proof of Lemma 3.1.
Among the two interior edges of $B^{1}$, let $f$ be a shortest one. Let $g\neq f$ be the other interior edge of $B^{1}$. Let $p$ be a shortest path in $\mathcal{S}(B)$. The relative interior $\mathring{p}$ of $p$ cannot intersect the relative interior of $f$ twice for those intersections would be crossings and $p$ and $f$ are both shortest paths because $B$ is good. So $\mathring{p}$ intersects $f$ less than four times. Then $\mathring{p}$ cannot intersect the relative interior of $g$ five times, for those intersections would be crossings, and $\mathring{p}$ would intersect $f$ in-between any two consecutive crossings with the relative interior of $g$. Altogether $p$ intersects $f$ and $g$ at most six times each. ∎
@@
Proposition 3.2 is similar to but different from a result of Löffler, Ophelders, Silveira, and Staals [22, Theorem 45] (building upon a ray shooting algorithm of Erickson and Nayyeri [10]), in which the authors provide an algorithm to transform a biface into a portalgon of bounded happiness, and of bounded combinatorial complexity. They extend their result from bifaces to portalgons $X$ such that the dual graph of $X^{1}$ in $\mathcal{S}(X)$ has at most one simple cycle, but unfortunately this does not include tubes. We extend their result to tubes to prove Proposition 3.2, reusing some of ideas developed in the core of the paper.
We need a few lemmas. The following is a corollary of [22, Theorem 45]:
@@
Proof.
By the result of Löffler, Ophelders, Silveira, and Staals [22, Theorem 45] we can compute in $O(1+\log h)$ time a portalgon $T$, whose surface is $\mathcal{S}(B)$, whose happiness is $O(1)$, and whose 1-skeleton $T^{1}$ has $O(1)$ edges. Without loss of generality the two vertices $b_{0}$ and $b_{1}$ of $B^{1}$ are also vertices of $T^{1}$, and we know which vertices of the polygons of $T$ correspond to $b_{0}$ and $b_{1}$.
We now describe how to compute, in constant time, from $T$, a good biface of $\mathcal{S}(T)$. The key thing is that we can exploit the fact that $T$ has $O(1)$ combinatorial complexity and happiness to compute by exhaustive search. First compute, in constant time, by exhaustive search, a shortest path $q$ between $b_{0}$ and $b_{1}$ in $\mathcal{S}(T)$: represent $q$ by its pre-image in the polygons of $T$. Then cut the polygons of $T$ along the pre-image of $q$: every time a polygon is cut in two along a segment $a$, the two edges issued of $a$ are not matched in the resulting portalgon (the goal is to cut the surface of $T$, not just changing $T$). Consider the resulting portalgon $D$. Then $\mathcal{S}(D)$ is homeomorphic to a closed disk. The two endpoints $b_{0}$ and $b_{1}$ of $q$ become a set $V$ of four vertices of $D^{1}$ that lie on the boundary of $\mathcal{S}(D)$. Every singularity of $\mathcal{S}(D)$ lies on the boundary of $\mathcal{S}(D)$ and belongs to $V$. Now replace $D$ by a triangular portalgon $D^{\prime}$, of the same surface, and such that the vertex set of $D^{\prime 1}$ is exactly $V$. This can be done for example by iteratively inserting vertex-to-vertex arcs in the faces of $D^{1}$ to make $D^{1}$ a triangulation, and by deleting a vertex $v$ of $D^{1}$ and its incident edges. When $v$ lies on the boundary of $\mathcal{S}(D)$, only the edges whose relative interior is included in the interior of $\mathcal{S}(D)$ are deleted. In the end, identify back the occurrences of $q$ on the boundary of $\mathcal{S}(D^{\prime})$, by matching the two corresponding sides of polygons in $D^{\prime}$, thereby obtaining a biface $B^{\prime}$ of $\mathcal{S}(B)$ such that $q$ is an interior edge of $B^{\prime}$. Change the other interior edge of $B^{\prime}$ if necessary so that $B^{\prime}$ is good. ∎
@@
Using the claim immediately the intersection of $p$ and $e$ has $O(1)$ connected components, so $p$ writes as a concatenation of $k=O(1)$ paths $p_{1},\dots,p_{k}$ such that for every $1\leq i\leq k$ the path $p_{i}$ is either included in $e$ or its relative interior is disjoint from $e$. Every edge $f\neq e$ of $T^{1}$ intersects $p_{i}$ less than 7 times: if $f$ is included in the boundary of $\mathcal{S}(T)$ then $f$ intersects $p_{i}$ at most once, otherwise Lemma 3.1 applies. So $f$ intersects $p$ less than $O(1)$ times. We proved that the segment-happiness of $T$ is $O(1)$. Then the happiness of $T$ is also $O(1)$ because the polygons of $T$ are all triangles. So we can compute a good biface whose surface is that of $T$ in constant time, exactly as in the proof of Lemma A.1. ∎
@@
Proof.
We may assume without loss of generality that no vertex of $Y$ in the interior of $S$ is incident to a loop edge, by cutting $S$ open at an interior loop edge and recursing on the resulting two triangulations otherwise. Euler’s formula gives $m-m_{1}+m_{2}=0$, where $m$, $m_{1}$, and $m_{2}$ count respectively the vertices, edges, and faces of $Y^{1}$. Double counting gives $3m_{2}=2m_{1}-2$ and $\sum_{v}\deg v=2m_{1}$, where the sum is over the vertices $v$ of $Y$. Then $\sum_{v}(6-\deg v)=4$. The two vertices of $Y$ on the boundary of $S$ have degree greater than or equal to four. So in the interior of $S$ every vertex of degree greater than ten must be compensated by a vertex of degree smaller than or equal to ten. ∎
@@
Let us now prove the claim. Consider the number $m^{\prime}$ of vertices of $X^{1}$ not incident to any loop edge that lie in the interior of $\mathcal{S}(X)$. By Lemma A.3, if $m^{\prime}>0$ before an iteration of the algorithm, then at least $m^{\prime}/2$ such vertices have degree smaller than or equal to ten. So $V$ contains at least $m^{\prime}/14$ vertices, which are deleted. Every non-deleted vertex that was incident to a loop edge before the iteration remains incident to a loop edge after the iteration. We proved that $m^{\prime}$ is divided by at least $14/13$ during the iteration, which proves the claim. ∎
Proof of Proposition 3.2.
Apply Lemma A.4, and replace $X$ in $O(n\log n)$ time by a concatenation of less than $3n$ bifaces whose edges are smaller than $(3n)^{c}L$ for some constant $c>0$. Each biface $B$ has segment-happiness $O(1+(3n)^{c}L/s)$; indeed the systole of $\mathcal{S}(B)$ is greater than or equal to the systole of $X$, so every segment $e$ in $\mathcal{S}(B)$ satisfies $h_{\mathcal{S}(B)}(e)=O(1+\ell(e)/s)$. Replace $B$ by a good biface whose surface is that of $B$ in $O(\log(n)+\log(2+L/s))$ time with Lemma A.1. Doing so for all bifaces takes $O(n\cdot(\log(n)+\log(2+L/s)))$ time in total. We crudely bound this running time from above by $O(n\log(n)\cdot\log(2+L/s))$. In the end apply Lemma A.2 repeatedly to merge those $O(n)$ good bifaces into a single good biface, in $O(n)$ total time. ∎